40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 4 12:46:27 2018, gautam, Update, ALS, Fiber ALS assay FiberALS.pdfFiberALS.svg.zip
    Reply  Thu Jan 11 14:22:40 2018, gautam, Update, ALS, Fiber ALS assay 
       Reply  Thu Jan 11 20:51:20 2018, gautam, Update, ALS, Fiber ALS assay IMG_5112.JPG
          Reply  Tue Jan 16 21:50:53 2018, gautam, Update, ALS, Fiber ALS assay 
             Reply  Thu Jan 18 00:35:00 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180118.pdfC2B4C1DD-6528-4067-9C13-6BD248629AD6.MOV
                Reply  Fri Jan 19 11:34:21 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180119.pdf
                   Reply  Fri Jan 19 23:04:11 2018, gautam, Update, ALS, Fiber ALS assay 
                      Reply  Wed Jan 24 00:33:31 2018, gautam, Update, ALS, Fiber ALS assay IMG_6866.JPG
                         Reply  Wed Jan 24 10:45:14 2018, gautam, Update, ALS, Fiber ALS assay 
                         Reply  Thu Jan 25 23:59:14 2018, gautam, Update, ALS, Fiber ALS assay 
                            Reply  Fri Jan 26 20:03:09 2018, gautam, Update, ALS, Fiber ALS assay 
                               Reply  Wed Jan 31 15:45:22 2018, gautam, Update, ALS, Fiber ALS assay IMG_6875.JPG
                                  Reply  Wed Jan 31 16:29:42 2018, gautam, Update, ALS, Modulation depths PMCmodDepth.pdfXPDH.pdf
                                  Reply  Wed Jan 31 16:33:53 2018, gautam, Update, ALS, ALS electronics at LSC rack 
                                     Reply  Wed Jan 31 22:32:11 2018, gautam, Update, ALS, ALS signal chain + power budget FiberALS_PowerBudget.pdfpreampProposed.pdf
                                        Reply  Thu Feb 1 01:24:56 2018, gautam, Update, ALS, D0902745 revamp underway PSLinterlock.png
                                           Reply  Fri Feb 2 00:26:34 2018, gautam, Update, ALS, D0902745 revamp underway 
                                              Reply  Fri Feb 2 13:16:55 2018, gautam, Update, ALS, ALS signals whitening switching ALS_whitening_switching.pdfALS_whitening_switching_works.pdf
                                                 Reply  Mon Feb 5 14:11:01 2018, gautam, Update, ALS, Huge harmonics in ALS channels ALSpowerlineHarmonics.pdf2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
                                                    Reply  Mon Feb 5 22:57:28 2018, gautam, Update, ALS, Huge harmonics in ALS channels 
                                                       Reply  Tue Feb 6 11:13:26 2018, gautam, Update, ALS, Possible source of ground loop identified A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
                                                          Reply  Tue Feb 6 22:55:51 2018, gautam, Update, ALS, Possible source of ground loop identified 
                                                       Reply  Thu Feb 8 00:33:20 2018, gautam, Update, ALS, D990694 characterization / THD measurement plan D990694THD_trial.pdf
                                                          Reply  Thu Feb 8 01:27:16 2018, Koji, Update, ALS, D990694 characterization / THD measurement plan 
                                                          Reply  Fri Feb 9 13:37:44 2018, gautam, Update, ALS, THD measurement trial THD_trial.pdf
                                                       Reply  Thu Feb 8 12:00:09 2018, gautam, Update, ALS, D990694 is NOT differential receiving 
                                                          Reply  Thu Feb 8 13:13:14 2018, gautam, Update, ALS, D990694 pulled out IMG_5131.JPGIMG_5133.JPGD990694-B_cuttingPlan.pdf
                                                             Reply  Thu Feb 8 18:10:36 2018, gautam, Update, ALS, D990694 pulled out 
                                              Reply  Wed Feb 7 15:51:15 2018, gautam, Update, ALS, D0902745 revamp complete 6613CD37-5014-44AE-B1FE-6F6A8137DF62.jpeg
                                                 Reply  Fri Feb 16 01:34:40 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180216.pdfIMG_5134.JPG
                                                    Reply  Tue Feb 20 23:08:27 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180220_diffOut.pdf
                                                       Reply  Thu Feb 22 00:09:11 2018, gautam, Update, ALS, D0902745 in-situ testing daughterBoard_TF.pdfdaughterBoard_noise.pdfIMG_6906.JPG
                                                          Reply  Sun Feb 25 00:03:12 2018, gautam, Update, ALS, Daughter board prototyping schematic.pdfdaughterBoard_TF.pdfdaughterBoard_noise.pdfTEK00000.PNGdaughterBoard_noise.pdf
                                                             Reply  Mon Feb 26 20:55:56 2018, rana, Update, ALS, Daughter board prototyping 
                                                                Reply  Tue Feb 27 21:10:45 2018, gautam, Update, ALS, Daughter board testing IMG_6912.JPGIMG_6913.JPGIMG_6914.JPG
                                                                   Reply  Wed Feb 28 19:13:25 2018, gautam, Update, ALS, ADC test for differential receiving in c1lsc c1lsc_ADCtest.png
                                                                   Reply  Fri Mar 2 01:45:06 2018, gautam, Update, ALS, new look ALS electronics IMG_6916.JPG
                                                                      Reply  Mon Mar 5 17:27:34 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                         Reply  Thu Mar 8 00:40:25 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                            Reply  Thu Mar 8 19:38:37 2018, gautam, Update, ALS, digital unwhitening of daughter board TFfit.pdf
                                                                               Reply  Thu Mar 8 23:50:27 2018, gautam, Update, ALS, First look at new ALS electronics BeatMouth_OOL.pdf
                                                                                  Reply  Fri Mar 9 01:07:01 2018, gautam, Update, ALS, First look at new ALS electronics ALS_vs_POXnoise.pdf
                                                                                     Reply  Mon Mar 12 22:08:31 2018, gautam, Update, ALS, Noisy POX NoisyPOX_20180312.pdf
                                                                                        Reply  Mon Mar 12 23:57:31 2018, gautam, Update, ALS, Noisy POX NoisyPOXandPOY_20180312.pdfnoisyPOX_Jan2018.pdf
                                                                                           Reply  Mon Mar 19 15:02:29 2018, gautam, Update, ALS, Noisy MC sensing 
                                        Reply  Thu Feb 1 15:31:12 2018, gautam, Update, ALS, ALS signal chain + power budget 
                                        Reply  Wed Feb 7 10:16:26 2018, gautam, Update, ALS, ALS signal chain + power budget 
                      Reply  Thu Jan 25 13:18:41 2018, gautam, Update, ALS, Fiber ALS assay preampProposed.pdf
Message ID: 13557     Entry time: Thu Jan 18 00:35:00 2018     In reply to: 13552     Reply to this: 13559
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS assay 

Summary:

I am facing two problems:

  1. The X arm beat seems to be broadband noisier than the Y arm beat - see Attachment #1. The Y-axis calibration is uncertain, but at least the Y beat has the same profile as the reference traces, which are for the green beat from a time when we had ALS running. There is also a rather huge ~5kHz peak, which I confirmed isn't present in the PDH error/control signal spectra (with SR785).
  2. The Y-arm beat amplitude, at times, "breathes" in amplitude (as judged by control room analyzer). Attachment #2 is a time-lapse of this behaviour (left beat is X arm beat, right peak is the Y arm peak) - I caught only part of it, the the beat note basically vanishes into the control room noise floor and then comes back up to almost the same level as the X beat. The scale is 10dB/div. During this time, the green (and IR for that matter) stay stably locked to the arm - you'll have to take my word for it as I have no way to sync my video with StripTool Traces, but I was watching the DC transmission levels the whole time. The whole process happens over a few (1< \tau <5) minutes - I didn't time it exactly. I can't really say this behaviour is periodic either - after the level comes back up, it sometimes stays at a given level almost indefinitely.

More details:

  • Spent some time today trying to figure out losses in various parts of the signal chain, to make sure I wasn't in danger of saturating RF amplifiers. Cabling from PSL table -> LSC rack results in ~2dB loss.
  • I will upload the updated schematic of the Beat-Mouth based ALS - I didn't get a chance to re-measure the optical powers into the Beat Mouth, as someone had left the Fiber Power Meter unplugged, and it had lost all of its charge angry.
  • The Demod boards have a nice "RF/LO power monitor" available at the backplane of the chassis - we should hook these channels up to the DAQ for long term monitoring.
    • The schematic claims "120mV/dBm" into 50ohms at these monitoring pins.
    • I measured the signal levels with a DMM (Teed with 50ohm), but couldn't really make the numbers jive - converting the measured backplane voltage into dBm of input power gives me an inferred power level that is ~5dBm higher than the actual measured power levels (measured with Agilent analyzer in Spectrum Analyzer mode).
  • Looking at the time series of the ALS I and Q inputs, the signals are large, but we are well clear of saturating our 16-bit ADCs.
  • In the brief periods when both beats were stable in amplitude (as judged by control room analyzer), the output of the Q quadrature of the phase tracker servo was ~12,000 cts - the number I am familiar with for the green days is ~2000cts - so naively, I would say we have ~6x the RF beat power from the Beat Mouth compared to green ALS.
  • I didn't characterize the conversion efficiency of the demod boards so I don't have a V (IF)/V (RF) number at the moment.
  • I confirmed that the various peaks seen in the X arm beat spectrum aren't seen in the control signal of the EX Green PDH, by looking at the spectrum on an SR785 (it is also supposedly recorded in the DAQ system, but I can't find the channel and the cable is labelled "GCX-PZT_OUT", which doesn't match any of our current channels).
    Note to self from the future: the relevant channels are: C1:ALS-X_ERR_MON_IN1 (green PDH error signal with x10 gain from an SR560) and C1:ALS-X_SLOW_SERVO_IN1 (green PDH control signal from monitor point - I believe this is DC coupled as this is the error signal to the slow EX laser PZT temp control). I've changed the cable labels at the X end to reflect this reality. At some point I will calibrate these to Hz.
  • The control room analyzer signals come from the "RF mon" outputs on the demod board, which supposedly couple the RF input with gain of -23dBm. These are then routed reverse through a power splitter to combine the X and Y signals, which is then plugged into the HP analyzer. The problem is not local to this path, as during the "breathing" of the Y beat RF amplitude, I can see the Q output of the phase tracker also breathing.

Next steps (that I can think of, ideas welcome!):

  1. For Problem #1 - usual debugging tactic of switching X and Y electronics paths to see if the problem lies in the light or in the electronics. If it is in the electronics, we can swap around at various points in the signal chain to try and isolate the problematic component.
  2. For Problem #2 - hook up the backplane monitor channels to monitor RF amplitudes over time and see if the drifts are correlated with other channels.
  3. There is evidence of some acoustic peaks, which are possibly originating from the fibers - need to track these down, but I think for a first pass to try and get the red ALS going, we shouldn't be bothered by these.

 

 

Attachment 1: IR_ALS_20180118.pdf  72 kB  | Hide | Hide all
IR_ALS_20180118.pdf
Attachment 2: C2B4C1DD-6528-4067-9C13-6BD248629AD6.MOV  4.134 MB
ELOG V3.1.3-